OpenStudy (anonymous):

Can someone help me find lim of x approaches 0 of (1-cos5x)/7x^2 ?

OpenStudy (anonymous):

I need to get there without l'hopital's rule, as we haven't studied that yet. I'm having problems with the algebraic manipulation.

OpenStudy (anonymous):

1-cos5x=1-(1-2sin^2(5x/2))=2sin^2(5x/2).......lim..[2sin^2(5x/2)]/7x^2=25/14

OpenStudy (anonymous):

Awesome, thanks!